Project

General

Profile

2017-09-08-EB

MICE Executive Board

8th September 2017; 15:00 BST

Phone details

Documents prepared for agenda items:

Operations Summary

  • Cycle 2017/02 starts on Sep 19 and goes to Oct 27
  • Goals:
    • take data with LH2 absorber
    • take data with SSD M2 powered
  • Broad brush run plan has been produced. Detail is now being filled in
  • Experimental shifts are fully allocated
  • BLOC shifts are not yet
  • Training plan has been produced
  • Start-up activities have begin
    • Yesterday we did a series of pulser runs to test detectors, RunControl, DAQ and monitoring
      • This went very well - the biggest problem is that the EMR could not be armed by the DAQ. This is
        an old problem, traceable to a readout board. Unfortunately we have no spares of this board. Francois is
        helping diagnose this problem today.
      • Tracker is being read out well, but we have many lost channels after the last warm-up. Melissa/Paul K are
        trying to see if we can live with this situation or if a channel reconfiguration can help. More next week.
      • Sep 14th : time allocated by ISIS during the machine physics run to do beam bump tuning and data-taking with beam. Caveat is that a lot of things are happening on the 14th so co-ordination and strict activity management will be critical to prevent the chaos.

Back!

Back

File list

2017-09-08-EB-Agenda.pdf (76.6 KB) Long, Kenneth, 06 September 2017 15:23

NuFactTalks.pdf (453 KB) Rogers, Chris, 08 September 2017 14:56